# Multi-Level Damage-Aware Graph Learning for Resilient UAV Swarm Networks

### v2

@@ IV-A Overall Framework of ML-DAGL
The proposed algorithm first constructs an input graph $\mathcal{G}_{in}(t_{0})=\{\mathcal{U}_{in},\mathcal{E}_{in}\}$ to resort the index of UAVs based on $\mathcal{I}_{R}$ and $\mathcal{I}_{D}$, i.e., $\mathcal{I}_{in}=\{r_{1},...,r_{N_{R}},d_{1},...,d_{N_{D}}\}$. Then, we can determine the input feature matrix as the position matrix of $\mathcal{G}_{in}$, that is
@@
The adjacent matrix of $\mathcal{G}_{in}$ is calculated as $\bm{A}_{in}=(a_{in,ij})$ based on $\mathcal{E}_{in}$, where $a_{in,ij}=1$ if and only if $u_{in,i}$ and $u_{in,j}$ form a communication link at time $t_{0}$.
The MBDA module transforms the input adjacent matrix $\bm{A}_{in}$ into several mDAGs via parallel branches. The $k$-th branch outputs an mDAG with the adjacent matrix $\bm{A}_{dag}^{k}$, which establishes links between remaining and damaged nodes within the range of $k$-hop neighbors. With the number of branches $K$, the module finally forms $K$ different mDAGs $\{\bm{A}_{dag}^{1},\bm{A}_{dag}^{2},...,\bm{A}_{dag}^{K}\}$.
The DGCN then outputs an optimal target position matrix $\bm{\hat{X}}_{gcn}^{k^{*}}$. Consequently, we can determine the optimal target position matrix for (P1) as
$\bm{\hat{X}}_{R}=[\bm{I}_{N_{R}}\enspace\bm{0}_{N_{D}}]\bm{\hat{X}}_{gcn}^{k^{*}}=[\bm{\hat{p}}_{r_{1}}^{k^{*}},\bm{\hat{p}}_{r_{2}}^{k^{*}},...,\bm{\hat{p}}_{r_{N_{R}}}^{k^{*}}]^{\top},$ (11)
@@
Notice that the proposed ML-DAGL is an optimization-based resilient algorithm; it requires online iterations to generate an optimal solution for the specific case with distinct input graph $\mathcal{G}_{in}$. To reduce the epoch of online iterations, we utilize the pre-trained model to initialize the parameters and skip the warm-up stage in training. Moreover, note that the shape of the input graph $\mathcal{G}_{in}$ only responds to the number of UAVs $N$ and parallel branches $K$. As a result, the pre-trained model can be adapted to different damage cases under varying $N_{D}$. In summary, the ML-DAGL requires limited online iterations and storage requirements, hence it is feasible for practical deployment on UAVs.
@@ IV-B Multi-Branch Damage Attention
The destroyed neighbors of the remaining nodes can effectively depict the damage model within local areas and guide the remaining nodes in filling the gaps in the damaged regions. To achieve this, we propose an MBDA module that extends the receptive fields via multi-hop dilation and applies damage attention to form the mDAG. The module structure is illustrated in Fig.2.
@@ IV-B1 Multi-hop dilation branches
Graph dilation aims to extend the receptive fields of nodes to their multi-hop neighbors. However, a static dilation size has limited performance when facing dynamic damage scales, since a large dilation size benefits the massive damage scenarios, while small-scale damage only requires a small dilation. In our approach, we develop $K$ parallel dilation branches to provide mDAG with multi-level damage-aware regions.
@@
$a_{in,ij}^{k}=\begin{cases}1,\quad&H_{in,ij}\leq k;\\ 0,&otherwise,\end{cases}$ (12)
where $H_{in,ij}$ denotes the number of hops between nodes $u_{in,i}$ and $u_{in,j}$ in $\mathcal{G}_{in}$. In other words, with the dilation size as $k$, the $\bm{A}_{in}^{k}$ will add dilated links between nodes whose number of hops is no larger than $k$. After the multi-hop dilation branches, we can get a series of dilated adjacent matrices $\{\bm{A}_{in}^{1},\bm{A}_{in}^{2},...,\bm{A}_{in}^{K}\}$.
@@ IV-B2 Damage attentive graph
where $\bm{A}_{in,r_{i}d_{j}}^{k}$ denotes the $k$-hop dilated links between $u_{r_{i}}$ and $u_{d_{j}}$, and $\odot$ denote the Hadamard product operator of matrices. Since the nodes in $\bm{A}_{in,r_{i}d_{j}}^{k}$ can be divided into the remaining-node and destroyed-node sets and no links exist between nodes within the same set, the mDAG is so-called a bipartite graph. This property splits the relationships between nodes distributed in uneven sub-nets, ensuring that the remaining nodes in the graph do not interact. After the damage attention processing, we have the mDAG sequence with $\{\bm{A}_{dag}^{1},\bm{A}_{dag}^{2},...,\bm{A}_{dag}^{K}\}$.
The parallel structure of MBDA module eliminates the need to pre-define a static dilation size $k$, which allows the algorithm to determine the optimal $k^{*}$ for each damage case. While available computational resources constrain the choice of the number of branches $K$, later analysis will demonstrate that the parallel structure significantly reduces both pre-training overhead and storage requirements.
@@ IV-C Dilated Graph Convolution Network
Building on the feature matrix $\bm{X}_{in}$ and the mDAG adjacent matrix sequence $\{\bm{A}_{dag}^{1},\bm{A}_{dag}^{2},...,\bm{A}_{dag}^{K}\}$ provided by MBDA, we present the structure of DGCN in Fig.3. The backbone of DGCN comprises an input normalization layer, the bipartite GCO layers, the nonlinear activation layers, and an output upscale layer.
@@ IV-C1 Structure of DGCN
Since each $\bm{A}_{dag}^{k}$ has a different dilation size $k$, we merge these adjacent matrices into a sparse block-diagonal matrix as the input for DGCN, i.e.,
$\bm{\dot{A}}_{dag}={\rm diag}(\bm{A}_{dag}^{1},\bm{A}_{dag}^{2},...,\bm{A}_{dag}^{K})\in\mathbf{S}^{KN}.$ (15)
@@
$\hat{\bm{\dot{X}}}_{gcn}={\rm GCN}(\bm{\dot{X}}_{in},\bm{\dot{A}}_{dag})=[\bm{\hat{X}}_{gcn}^{1\top},\bm{\hat{X}}_{gcn}^{2\top},...,\bm{\hat{X}}_{gcn}^{K\top}]^{\top},$ (17)
where each $\bm{\hat{X}}_{gcn}^{k}=[\bm{\hat{p}}_{r_{1}}^{k},...,\bm{\hat{p}}_{r_{N_{R}}}^{k},\bm{\hat{p}}_{d_{1}}^{k},...,\bm{\hat{p}}_{d_{N_{D}}}^{k}]^{\top}\in\mathbb{R}^{N\times 2}$ is a possible solution for (P1). Among these $K$ possible solutions, we can calculate the recovery time of the $k$-th solution as
$T_{rc}^{k}=\mathop{\rm max}\limits_{r_{i}\in\mathcal{I}_{R}}\frac{\|\bm{\hat{p}}_{r_{i}}^{k}-\bm{p}_{r_{i}}(t_{0})\|}{v_{max}}.$ (18)
@@ IV-C2 Bipartite GCO
Subsequently, we extend the $g_{\theta}$ into a trainable B-GCO layer. In each B-GCO layer, we introduce a weight matrix for feature linear transformation, which serves as the trainable parameter. Specifically, in the $q$-th B-GCO layer, the feature matrix is processed by the GCO as
@@ IV-C3 DGCN backbone
As shown in the right part of Fig.3, a position normalization layer takes $\bm{\dot{X}}_{in}$ as input and down-scales the features into the range of $(-1,1)$. Let $\bm{p}_{c}=\frac{1}{N}\sum_{i=1}^{N}\bm{p}_{i}(t_{0})$ denote the center point of the USNET at time $t_{0}$, hence we can define the scale ratio $D=\mathop{\rm{max}}\limits_{i\in\mathcal{I}}\|\bm{p}_{c}-\bm{p}_{i}(t_{0})\|$ is the maximum distance between nodes and $\bm{p}_{c}$. The normalized feature matrix as be calculated as
@@
After $L$ residual blocks, the final B-GCO layer brings the feature dimension back to 2 with the weight matrix set as a $d_{s}$by-2 matrix. In addition, we introduce a tanh function to map the feature into the range of $(-1,1)$. Therefore, the DGCN has a total of $Q=2L+2$ GCO layers with the output
@@ IV-C4 Loss function design
Since the output of DGCN $\hat{\bm{\dot{X}}}_{gcn}$ contains $K$ possible solutions according to (17), we can calculate a loss for each solution $\bm{\hat{X}}_{gcn}^{k}$ respectively. To transform the equality constraint (9) as a penalty component in the loss function, we rewrite the problem (P1) as
[…]
@@ IV-D Theoretical Analysis of ML-DAGL
To analyze ML-DAGL, we need to address whether the solution can guarantee connectivity recovery and assess its performance under worst-case scenarios. We first demonstrate that the B-GCO converges for the feature matrix of the mDAG. Subsequently, we derive the theoretical upper bound for the recovery time, relying on the final convergent matrix of the B-GCO.
@@ IV-D1 Convergence of B-GCO
Since the $\bm{\dot{A}}_{dag}$ is block-diagonal, we can regard the B-GCO on feature matrix $\bm{\dot{X}}_{in}$ as applying B-GCO on $\bm{X}_{in}$ for $K$ times with different $\bm{A}_{dag}^{k}$. For each B-GCO with $\bm{A}_{dag}^{k}$, we can prove that it is a contraction operation for each iteration.
@@
Especially, when the B-GCO takes $\bm{X}_{in}$ as the input, $\bm{\bar{X}}_{gco}\equiv[\bm{p}_{c},\bm{p}_{c},...,\bm{p}_{c}]^{\top}$ holds when the $k$-hop mDAG is connected.
@@
Proposition 1 demonstrates that, after a sufficient number of iterations, the B-GCO output $\bm{\hat{X}}_{gco}^{k}$ converges to $\bm{\bar{X}}_{gco}^{k}$. Since the remaining graph constructed from $\bm{\bar{X}}_{gco}\equiv[\bm{p}_{c},\bm{p}_{c},...,\bm{p}_{c}]^{\top}$ is fully connected, we need to find the number of parallel branches $K$ that ensure at least one connected mDAG with $k\leq K$, and determine a proper $\epsilon$ to meet the convergence constraint for each mDAG with different dilation size $k$.
@@ IV-D2 Choice of hyper-parameters
As the number of nodes $N$ and the extent of damage $N_{D}$ increase, a larger dilation size $k$ is required to ensure that the mDAG remains connected. However, the number of parallel branches $K$ limits the maximum value of $k$ with $k_{max}=K$. It can be easily observed that if a $k$-hop mDAG with $\bm{A}_{dag}^{k}$ is connected, then for every $k_{i}<k\leq K$, the $k_{i}$-hop mDAG is connected. Therefore, we only need to ensure that the mDAG is connected in the worst-case scenario. Denote $H_{max}$ as the maximum hop number in the original USNET, we set the choice of $K$ as
$K=\lfloor\frac{H_{max}+1}{2}\rfloor.$ (35)
[…]
For the choice of $\epsilon$, note that $\|\bm{A}_{dag}^{k}\|_{\infty}$ is positively correlated to the dilation size $k$, but with a limit of $\|\bm{A}_{dag}^{k}\|_{\infty}<N$. To ensure that the convolution kernel $g_{\theta}$ remains isomorphic for different $k$ during the batch process, we set $\epsilon=\frac{1}{N}<\frac{1}{\|\bm{A}_{dag}^{k}\|_{\infty}}$ for the bipartite graph convolution to meet the convergence constraint. With the $K$ in (35), the ML-DAGL algorithm must provide a solution that guarantees connectivity restoration, hence the convergence holds.
@@ IV-D3 Upper bound of recovery time
Since the final solution $\bm{\hat{X}}_{gcn}^{k^{*}}$ will approach $\bm{\bar{X}}_{gcn}$ when the mDAG is connected, the solution under the worst case refers $\bm{\hat{X}}_{R}=[\bm{p}_{c},...,\bm{p}_{c}]^{\top}$ that guarantees the restoration of connectivity. Therefore, we can give an upper bound for the recovery time as
$T_{rc}^{max}=sup\{T_{rc}^{k^{*}}\}=\frac{\mathop{\rm max}\limits_{r_{j}\in\mathcal{I}_{R}}\|\bm{p}_{c}-\bm{p}_{r_{j}}(t_{0})\|}{v_{max}}.$ (36)
In practical applications, the model often does not need to iterate to $\bm{\bar{X}}_{gco}$ to find a suitable solution, making early termination essential. According to Appendix A, when the number of iterations $q$ is small, the location of each remaining UAV tends to aggregate around the local damaged region. This characteristic contributes to the performance advantage of the ML-DAGL algorithm.

### v3

@@ IV-A Overall Framework of ML-DAGL
The proposed algorithm first constructs an input graph $\mathcal{G}_{in}(t_{0})=\{\mathcal{U}_{in},\mathcal{E}_{in}\}$ to resort the order of UAVs based on node status, i.e., $\mathcal{U}_{in}=\{u_{r_{1}},...,u_{r_{N_{R}}},u_{d_{1}},...,u_{d_{N_{D}}}\}$. Then, we can determine the input feature matrix as the position matrix of $\mathcal{G}_{in}$, that is
@@
The adjacent matrix of $\mathcal{G}_{in}$ is calculated as $\bm{A}_{in}=(a_{in,ij})$ based on $\mathcal{E}_{in}$, where $a_{in,ij}=1$ if and only if link $e_{in,ij}$ exists.
The MBDA module transforms the input adjacent matrix $\bm{A}_{in}$ into several mDAGs via parallel branches. The $k$-th branch outputs an mDAG with its adjacent matrix $\bm{A}_{dag}^{k}$, which establishes links between remaining and damaged nodes within the range of $k$-hop neighborhood. With the number of branches $K$, the module finally forms $K$ different mDAGs $\{\bm{A}_{dag}^{1},\bm{A}_{dag}^{2},...,\bm{A}_{dag}^{K}\}$.
Then, DGCN applies convolution on the series of mDAGs and outputs a target position matrix $\bm{\hat{X}}_{gcn}^{k^{*}}$. Consequently, we can determine the optimal target position matrix for (P1) as
$\bm{\hat{X}}_{R}=[\bm{I}_{N_{R}}\kern 5.0pt\bm{0}_{N_{D}}]\bm{\hat{X}}_{gcn}^{k^{*}}=[\bm{\hat{p}}_{r_{1}}^{k^{*}},\bm{\hat{p}}_{r_{2}}^{k^{*}},...,\bm{\hat{p}}_{r_{N_{R}}}^{k^{*}}]^{\top},$ (12)
@@
Since the proposed ML-DAGL is an optimization-based transductive algorithm, it requires additional online iterations to generate an optimal solution for a specific input graph $\mathcal{G}_{in}$. To reduce online iterations, pre-trained models are deployed to initialize the parameters of DGCN and skip the warm-up stage in training. Note that all mDAGs share an unified shape related to the number of UAVs $N$, a pre-trained model for a certain swarm scale can be adapted to distinct damage cases under varying $N_{D}$. In summary, the ML-DAGL requires limited online iterations and storage requirements, hence it is feasible for practical deployment on UAVs.
@@ IV-B Multi-Branch Damage Attention
The destroyed neighbors of the remaining nodes can effectively depict the damage model within local areas and guide the remaining nodes in filling gaps on the damaged regions. To achieve this, we propose the MBDA module that extends the receptive fields via multi-hop dilation and applies damage attention to form mDAGs. The module structure is illustrated in Fig.2.
@@ IV-B1 Multi-hop dilation branches
To enhance the sparse topology of the input graph $\mathcal{G}_{in}$, we add edges between nodes and their multi-hop neighbors for receptive fields dilation. However, a static dilation size has limited performance when facing dynamic damage scales, since a large dilation size benefits the massive damage scenarios, while small-scale damage only requires a small dilation. In our approach, we develop $K$ parallel dilation branches to provide mDAG with multi-level damage-aware regions.
@@
[…]
where $H_{in,ij}$ denotes the number of hops between nodes $u_{in,i}$ and $u_{in,j}$ in $\mathcal{G}_{in}$. In other words, with the dilation size as $k$, $\bm{A}_{in}^{k}$ will form links between nodes whose number of hops is no larger than $k$. After the multi-hop dilation branches, we can get a series of dilated adjacent matrices $\{\bm{A}_{in}^{1},\bm{A}_{in}^{2},...,\bm{A}_{in}^{K}\}$.
@@ IV-B2 Damage attentive graph
where $\bm{A}_{in,r_{i}d_{j}}^{k}$ denotes the $k$-hop dilated links between $u_{r_{i}}$ and $u_{d_{j}}$, and $\odot$ denote the Hadamard product operator of matrices. Since nodes in $\bm{A}_{in,r_{i}d_{j}}^{k}$ can be divided into the remaining-node and destroyed-node sets and no links exist between nodes within the same set, the mDAG is so-called a bipartite graph. This property splits the relationships between nodes distributed in uneven sub-nets, ensuring that the remaining nodes in the graph do not interact. After the damage attention processing, we have the mDAG sequence with $\{\bm{A}_{dag}^{1},\bm{A}_{dag}^{2},...,\bm{A}_{dag}^{K}\}$.
The parallel structure of MBDA module allows the algorithm to always determine an optimal $k^{*}$ for each damage case, adapting to varying scales of swarm $N$ and damage $N_{D}$. While available computational resources constrain the choice of the number of branches $K$, later analysis will demonstrate that the parallel structure significantly reduces both pre-training overhead and storage requirements.
@@ IV-C Dilated Graph Convolution Network
The mDAGs formed by MBDA module have special bipartite topologies, basic graph convolution operation (GCO) may lead to extensive over-smoothing problem after two GCO layers. In this paper, we propose the DGCN to apply bipartite GCO on mDAGs with residual blocks, achieving gradually contracting feature aggregation. Building on the feature matrix $\bm{X}_{in}$ and the mDAG adjacent matrix sequence $\{\bm{A}_{dag}^{1},\bm{A}_{dag}^{2},...,\bm{A}_{dag}^{K}\}$ provided by MBDA, we present the structure of DGCN in Fig.3.
@@ IV-C1 Structure of DGCN
For the series of mDAGs $\bm{A}_{dag}^{k}$ that have different dilation sizes, we merge these adjacent matrices into a sparse block-diagonal matrix as the input for DGCN, i.e.,
$\bm{\dot{A}}_{dag}={\rm diag}(\bm{A}_{dag}^{1},\bm{A}_{dag}^{2},...,\bm{A}_{dag}^{K})\in\mathbb{S}^{KN}.$ (16)
@@
$\hat{\bm{\dot{X}}}_{gcn}={\rm DGCN}(\bm{\dot{X}}_{in},\bm{\dot{A}}_{dag})=[\bm{\hat{X}}_{gcn}^{1\top},\bm{\hat{X}}_{gcn}^{2\top},...,\bm{\hat{X}}_{gcn}^{K\top}]^{\top},$ (18)
where each $\bm{\hat{X}}_{gcn}^{k}=[\bm{\hat{p}}_{r_{1}}^{k},...,\bm{\hat{p}}_{r_{N_{R}}}^{k},\bm{\hat{p}}_{d_{1}}^{k},...,\bm{\hat{p}}_{d_{N_{D}}}^{k}]^{\top}\in\mathbb{R}^{N\times 2}$ is a possible solution for (P1). The recovery time of the $k$-th solution is calculated as
$T_{rc}^{k}=\mathop{\rm max}\limits_{u_{r_{i}}\in\mathcal{U}_{R}}\frac{\|\bm{\hat{p}}_{r_{i}}^{k}-\bm{p}_{r_{i}}(t_{0})\|}{v_{max}}.$ (19)
@@ IV-C2 Bipartite GCO
Subsequently, we extend the $g_{\theta}$ into a trainable B-GCO layer. In each B-GCO layer, we introduce a weight matrix for feature linear transformation, which serves as the trainable parameter. Specifically, in the $q$-th B-GCO layer, the feature matrix is processed as
@@ IV-C3 DGCN backbone
As shown in the right part of Fig.3, DGCN backbone comprises an input normalization layer, the bipartite GCO layers, the nonlinear activation layers, and an output upscale layer.
[…]
@@
After $L$ residual blocks, the final B-GCO layer brings the feature dimension back to 2 with the weight matrix set as a $d_{s}\times 2$ dimensional matrix. In addition, we introduce a tanh function to map the feature into the range of (1,1)(-1,1). Therefore, the DGCN has a total of $Q=2L+2$ GCO layers with the output
@@ IV-C4 Loss function design
Note that the output of DGCN $\hat{\bm{\dot{X}}}_{gcn}$ contains $K$ possible solutions according to (18). Hence, we can calculate a loss $\mathcal{L}^{k}$ under each solution $\bm{\hat{X}}_{gcn}^{k}$ respectively, where the design of $\mathcal{L}^{k}$ should be consistent with (P1). Specifically, we rewrite the problem (P1) as
$({\rm P}1^{\dagger}):\quad\mathop{\rm min}\limits_{\bm{\hat{X}}_{gcn}^{k}}\quad$ $T_{rc}^{k}=\mathop{\rm max}\limits_{u_{r_{i}}\in\mathcal{U}_{R}}\frac{\|\bm{\hat{p}}_{r_{i}}^{k}-\bm{p}_{r_{i}}(t_{0})\|}{v_{max}}$ (31)
@@ IV-D Theoretical Analysis of ML-DAGL
To analyze ML-DAGL, we need to address whether the solution can guarantee connectivity recovery and assess its performance under worst-case scenarios. We first demonstrate that the B-GCO always converges to a feasible solution on a connected mDAG, and derive the selection of hyper-parameters to form a connected mDAG. Subsequently, we analyze the theoretical upper bound for the recovery time.
@@ IV-D1 Convergence of B-GCO
Since $\bm{\dot{A}}_{dag}$ is block-diagonal, we can regard the B-GCO on feature matrix $\bm{\dot{X}}_{in}$ as applying B-GCO on $\bm{X}_{in}$ for $K$ times with different $\bm{A}_{dag}^{k}$. For each B-GCO with $\bm{A}_{dag}^{k}$, we can prove that it is a contraction operation for each iteration as follows.
@@
Especially, when B-GCO takes $\bm{X}_{in}$ as the input, $\bm{\bar{X}}_{gco}\equiv[\bm{p}_{c},\bm{p}_{c},...,\bm{p}_{c}]^{\top}$ holds when the $k$-hop mDAG is connected.
@@
Proposition 1 demonstrates that, after a sufficient number of iterations, the output matrix $\bm{\hat{X}}_{gco}^{k}$ of a connected mDAG will converge to $\bm{\bar{X}}_{gco}^{k}$. Therefore, B-GCO can always provide a feasible solution for (P1) since the remaining graph constructed from $\bm{\bar{X}}_{gco}$ is fully connected. To guarantee the convergence of B-GCO on varying damage cases, we need to find a proper $K$, the number of parallel branches, to ensure that at least one mDAG is connected, and determine a proper $\epsilon$ to meet the convergence constraint.
@@ IV-D2 Choice of hyper-parameters
As the scales of nodes $N$ and damage $N_{D}$ increase, mDAG with a larger dilation size $k$ is more likely to be connected. It can be easily observed that once a $k$-hop mDAG with $\bm{A}_{dag}^{k}$ is connected, then for every $k<k_{i}\leq K$, the $k_{i}$-hop mDAG is connected. Note that $k$ is limited by $k\leq K$, therefore, we only need to ensure that the mDAG on the $K$-th branch is always connected in the worst-case scenario. Denote $H_{max}$ as the maximum number of hops in the original USNET, we set the choice of $K$ as
$K=\left\lfloor\frac{H_{max}+1}{2}\right\rfloor.$ (36)
where $\lfloor\cdot\rfloor$ denotes the floor function.
For the choice of $\epsilon$, we have $\|\bm{A}_{dag}^{k}\|_{\infty}$ positively correlated to the dilation size $k$. Note that constraint $\|\bm{A}_{dag}^{k}\|_{\infty}<N$ always holds. To ensure that the convolution kernel $g_{\theta}$ remains isomorphic for different $k$ during the batch process, we set $\epsilon=\frac{1}{N}<\frac{1}{\|\bm{A}_{dag}^{k}\|_{\infty}}$ for the bipartite graph convolution to meet the convergence constraint. With the $K$ in (36), the ML-DAGL algorithm must provide a solution that guarantees connectivity restoration, hence the convergence holds.
@@ IV-D3 Upper bound of recovery time
Since the final solution $\bm{\hat{X}}_{gcn}^{k^{*}}$ will approach $\bm{\bar{X}}_{gcn}$ when the mDAG is connected, the solution under the worst case refers to $\bm{\hat{X}}_{R}=\bm{\bar{X}}_{gco}=[\bm{p}_{c},...,\bm{p}_{c}]^{\top}$ that guarantees the restoration of connectivity. Therefore, we can give an upper bound for the recovery time as
$T_{rc}^{max}=sup\{T_{rc}^{k^{*}}\}=\frac{\mathop{\rm max}\limits_{u_{r_{j}}\in\mathcal{U}_{R}}\|\bm{p}_{c}-\bm{p}_{r_{j}}(t_{0})\|}{v_{max}}.$ (37)
In practical applications, the model often finds a more efficient solution based on early termination before it converges to $\bm{\bar{X}}_{gco}$. According to Appendix A, when the number of iterations is small, the remaining UAV tends to aggregate around their neighboring damaged region. This characteristic contributes to the performance advantage of the ML-DAGL algorithm.